solve: sqrt (r - 1) + 5 = 2
a) (10)
b) (8)
c) (50)
d) none or 0

Answer by checkley77(12844)

sqrt (r - 1) + 5 = 2
sqrt(r-1)=2-5
sqrt(r-1)=-3 square both sides.
r-1=9
r=9+1
r=10 ans. A)